OpenStudy (anonymous):

Using laws of logic, prove that (Q^ notR) ) =>notP is logically equivalent to (P ^ Q) ) => R

OpenStudy (anonymous):

i dont like truth tables =/

OpenStudy (jamesj):

(Q and not R) => not P = not( (Q and not R) and not (not P)) = not( (Q and not R) and P) = not ( (P and Q) and not R) = (P and Q) => R

OpenStudy (anonymous):

would it be a truth table though? i can do those no problem, but it says using laws of logic.... does that make a difference?

OpenStudy (jamesj):

The important result you need, and I've used, is that P => Q is equivalent to not (P and not Q)

OpenStudy (anonymous):

one way to show two logical statements are equivalent is to show their truth tables are equivalent. thats just one way to do it though. James' argument is a lot shorter.

OpenStudy (anonymous):

ok thanks very much guys!